What is the value of Luis Tiant's actual cap from the 1975 World Series Game 6?
This hat is extremely rare and obviously one of a kind. To the right person, it would fetch upwards of $20,000.
What does qualitative observation mean?
A Qualitative observation is an observation that is found using data without numbers. (hence the term quality instead of quantity which is quantitative)
Ex: the dog is brown
This orange has a peel
People har hair
Sometimes, simple examples help the most.
Take the 5 numbers, 1,2,3,4,5. The number in the middle is 3 so it is the median.
Now take 4 numbers, 1,2,3,4. There is no single middle integer. So we take the middle two, add them, and divide by 2. (2+3)/2=5/2 which is the median.
Note that when we have an even number of entries, we must also use this method, but we can just pick the middle number when there is an odd number of entries.
If all students in a class scored 100 on an exam what is the standard deviation?
Standard deviation is 0.
Two flips can have four possible results:
T - T
T - H
H - T
H - H
The question defines 'success' as either H-T or T-H.
Two successes out of four possibilities = probability of 2/4 = 1/2 = 50% .
8/100 or 2/25
There is no number. 100 is the only one that has a hundreds digit. It's impossible.
What is 23 over 150 as a percentage?
Expressed as a percentage, 23/150 x 100 = 15.3 recurring (that is, 15.3333..) percent.
How many 7 number combinations can be made with 8 numbers?
8*7*6*5*4*3*2
8P7
=40,320
1. If you are really asking for combinations, in other words, if the order of the numbers chosen doesn't matter :
If each number can be selected only once, this is the number of choices of 7 from 8 _without replacement_:
8 choose 7 = 8
If each number can be selected more than once, this is the number of choice of 7 from 8 _with replacement_:
( 8 + 7 - 1 ) choose 7 = 3432
2. If the order of the numbers chosen matters as well as the combinations :
If each number can be selected only once, then there are 8 choices for the 1st number, 7 for the next, 6 for the one after that, and so on, or 8*7*6*5*4*3*2, as in the first answer.
However, if each number can be selected more than once, then there are 8 choices for the 1st number, (still) 8 choices for the next, 8 choices for the one after that, and so on, or 8*8*8*8*8*8*8 = 2097152.
How many 5 digit numbers can be formed using the digits 1 2 3 4 5 6 7?
Without repeated digits, the answer is 7*6*5*4*3 = 2520
Does the y-coordinate in the ordered pair correspond to the input or the output?
Neither, as a graph can be read from either axis to get a value on the other axis.
However, the y coordinate is usually considered the dependent variable as an equation is usually written as y = f(x), so it could be considered the output.
What are the four digit figures possible?
You start with 0001 and count your way up to 9999.
There's a good reason that they're not listed here for you. There are quite
a goodly number of them. Let's figure out how many there are, shall we ?
.0001 to .9999 . . . . . 9,999 numbers
1.001 to 9.999 . . . . . 8,999 numbers
10.01 to 99.99 . . . . . 8,999 numbers
100.1 to 999.9 . . . . . 8,999 numbers
1000 to 9999 . . . . . . 9,000 numbers
Total . . . . . 45,996 numbers that require 4 digits
18 000 meters are
180 000 dm
---------------------------------------------------------------------------------------------
10 dm = 1 meter,
or, 1/10 meter = 1 dm
So 18 000 meters = 18 000 meter [10 dm/1 meter] = 180 000 dm
What does it mean for the probability of an event to be 100 percent or 1 explain in two paragraphs?
Paragraph #1:
A probability of 1, or 100%, indicates certainty. For example:
-- The probability of your room remaining full of air is virtually 100%.
-- The probability of the sun rising tomorrow is believed to be 100%.
-- An event must either occur or not occur, so if you take the probability
of it occurring, and add the probability that it won't occur, they must
add up to 100% .
Paragraph #2:
It was too much trouble for me to look up this information and then write about it.
So I posted the question on a Q&A website, and somebody there gave me the answer.
If it's wrong then it's his fault, but if it's correct, I'll be happy to take full credit for it.
Either way, I understand that even though I've given you an answer, you don't need it
and I haven't learned anything.